OpenStudy (anonymous):

Hector is preforming a chemistry experiment that requires 140 milliliters of a 30% copper sulfate solution. he has a 25% copper sulfate solution and a 60% copper sulfate solution. How many milliliters of each solution should he mic to obtain the needed solution?

OpenStudy (nottim):

There is a specific portion of this site for Chemistry. You might get better help over there.

OpenStudy (anonymous):

ok thanks! ill try that

OpenStudy (amistre64):

this is an algebra word problem

OpenStudy (anonymous):

yea there wasnt a chemistry section anyways.... :) thanks for trying tho

OpenStudy (nottim):

Yeah. Sorry bro/girl.

OpenStudy (nottim):

I am a dumb dumb.

OpenStudy (anonymous):

yea there wasnt a chemistry section anyways.... :) thanks for trying tho

OpenStudy (anonymous):

its ok you tried to help

OpenStudy (amistre64):

amount@25%=140(.30-.60)/(.25-.60) = 120 leaving 20 for the 60%

OpenStudy (anonymous):

yea there wasnt a chemistry section anyways.... :) thanks for trying tho

OpenStudy (anonymous):

ok ill try that

OpenStudy (amistre64):

i find the whole substitution equations and eliminations rather laborious, so I just go straght for it :)

OpenStudy (anonymous):

i love the one step method. no muss, no fuss!

OpenStudy (anonymous):

satellite73 how would you solve this?

OpenStudy (anonymous):

just wish i understood it better

OpenStudy (anonymous):

i would say "put amount of 60% copper solution as x, then since the total is 140 the amount of 25% solution must be 140 - x" then solve \[.6x+.25(140-x)=.3\times 140\] but i really do like the other way better

OpenStudy (anonymous):

ok thank you! :) everyone says im different and i like to learn algebra in a different way so this way is easyer for me... :)

OpenStudy (anonymous):

is that .3 times 140 or .3x times 140?

OpenStudy (amistre64):

a + b = 140 ; *-.60 a(.60) + b(.25) = 140(.30) -a(.60) +b(-.60) =140(-.60) a(.60) + b(.25) = 140(.30) --------------------------- b(.25-.60) = 140(.30-.60) b @ 25% = 140(.30-.60)/(.25-.60)